When writing the expression as one logarithm, how far do you go with this problem? log cubed squared root of x + log x^4 - log x^3?
v/r Dean

Answer by venugopalramana(3286)

log cubed squared root of x + log x^4 - log x^3?
IT IS ASSUMED THAT BY log cubed squared root of x YOU MEAN
LOG((X)^0.5)^3)...ASSUMING THAT AND USING FORMULA LOG (X)^N=N*LOG X,
AND ((X)^M)^N)=(X)^M*N ......
WE GET
log cubed squared root of x + log x^4 - log x^3 = LOG (X)^1.5+4*LOG X -3*LOG X
=(1.5+4-3)*LOG x = 2.5 LOG X=LOG (X)^2.5
ON THE OTHER HAND IF IT IS MEANT AS {LOG (X)^0.5}^3}...THEN IT WILL BE EQUAL TO
={0.5*LOG (X)}^3+4*LOG X -3*LOG X=0.125*{LOG X}^3+LOG X